# Supplement to “Steering LLMs Towards Unbiased Responses: A Causality-Guided Debiasing Framework”

### v1

@@ 4.1.1 Experimental Settings (WinoBias)
  • ICL with contrastive examples: provides contrastive examples followed by original question (Figure 2 middle). The examples are considered to be contrastive as the answer remains unchanged with different pronouns. In WinoBias, we use the same 16 ICL examples as in Si et al. (2022).
  • Zero-shot COT (Kojima et al., 2022): asks the model to also “think step by step” after the original question. Shaikh et al. (2022); Ganguli et al. (2023) design similar COT approaches to mitigate bias in LLMs.
Evaluation Metrics
We measure the performances of LLMs on pro and anti sentences in terms of accuracy, and the gap between the two indicates the level of gender bias exhibited by the models (a smaller gap is better).
Our Method: Reduce + Fact
We propose Reduce + Fact as one way of encouraging bias-free reasoning collectively with discouraging biased reasoning. As shown in Figure 2 right panel, we first create two gender-agnostic sentences by replacing the gender pronoun with the two occupations that appeared in the original sentence. Then we ask the model a factual question: which sentence is more likely to happen in real life, as this question does not contain any gender-related information. We then prepend its answer (to the factual question) before asking the original question. This allows us to distill LLM’s non-gender-related world knowledge (Strategy III) and nudge it to explicitly reason with this fact during coreference resolution (Strategy I). On top of encouraging bias-free reasoning (Strategies I and III), we also prompt that both occupations are equally likely to be male or female to counteract existing selection bias (Strategy II), discouraging biased reasoning.
One caveat of utilizing LLM’s world knowledge is that the performance of Reduce + Fact will increase as the capabilities of the LLM grow, since better world knowledge will further help LLM answer both factual and original questions.
4.1.2 Our Results (WinoBias)
We compare Reduce + Fact with the three baseline approaches on both coreference tasks and summarize our results in Table 3.2. We expect all LLMs to achieve higher accuracies on sentences containing pro-stereotypical pronouns (pro sentences) compared to these with anti-stereotypical pronouns (anti sentences) due to historical biases in the training data.
Coreference Resolution Requiring World Knowledge
For Type I tasks, world knowledge is required to perform coreference resolution. The Default prompting shows significant biases, with large gaps in accuracy between pro sentences and anti sentences for all models. The method Zero-shot COT marginally reduces the bias, as seen in the smaller gaps, but its performance is lower on both pro and anti scenarios for GPT-3 and GPT-3.5 when compared with Default, and it only marginally improves the general performance when applied to more capable ones (Claude 2 and GPT-4). For ICL with counterfactual examples, although the gap becomes larger for less capable models (GPT-3 and GPT-3.5), it can further reduce biases with more capable LLMs (Claude 2 and GPT-4), especially for GPT-4 (with a gap of 9.23%).
Remarkably, Reduce + Fact, which encourages bias-free reasoning while discouraging biased reasoning, substantially decreases the bias across all LLMs. The achieved minimal gaps, with GPT-4 exhibiting a mere 2.17% gap and 94.57% accuracy on anti sentences, demonstrate significantly greater effectiveness compared to other approaches.
Coreference Resolution Only with Syntactic Information
For Type II tasks, coreference can be resolved only with syntactic cues. All models generally show higher accuracies with smaller gaps. However, since there exists a shortcut in the Type II task – the correct answer is always the second entity in the sentence – a smaller gap may not necessarily indicate less bias. Still, Reduce + Fact outperforms the three baselines, particularly with GPT-4, achieving a near-negligible gap of 0.13%, suggesting that our method is highly effective in encouraging models to utilize non-gendered cues for coreference resolution.
Summary
These experimental results suggest prompt designs that encourage bias-free reasoning and discourage biased reasoning are effective at mitigating gender biases in LLMs by directing them to rely more on non-gender-related world knowledge (Strategies I and III) and less on gender shortcuts (Strategy II), thus promoting fairer and less biased responses. Also, the performance gap between pro sentences and anti sentences decreases as the LLMs become more capable, which may indicate that LLMs are less prone to assign occupations with stereotypical gender pronouns as their general (reasoning) capabilities grow.
Figure 5: Performance comparison on Discrim-Eval across three demographics. The bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Encouraging bias-free reasoning universally decreases the relative gap when added with methods that reduce biased reasoning.
4.1.3 Ablation studies
We further conduct ablation studies to individually examine the effectiveness of the two types of approaches, i.e., encouraging bias-free reasoning and discouraging biased reasoning. We divide the prompt Reduce + Fact into two parts: the Reduce Only part prompting both men and women can perform either occupation equally, and the Fact Only part telling the model which sentence(s) it regards to be more viable in real life.
Moreover, we categorize each LLM’s response into four groups: (i) TT: where the LLM answers both the factual question and the original question correctly, (ii) TF: where the LLM only answers the factual question correctly, (iii) FT: where the LLM only answers the original question correctly, and (iv) FF: where the LLM answers wrongly on both questions.
@@ 4.1.3 Ablation studies
Table 2: Error Analysis on WinoBias Type I coreference task. We divide the models’ responses into 4 categories to better understand the cause of their errors and success. $F$ denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge. TF denotes the coreference errors caused by gender bias as only factual questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the factual questions get wrong but original questions are correctly answered.
GPT-4   Reduce + Fact 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Fact Only 91.45 95.93 5.29 0.14 0.54 0.81 2.71 3.12 Reduce Only 70.69 88.20 25.78 8.28 0.95 1.76 2.58 1.76 Default 69.47 87.25 26.87 8.68 1.36 2.44 2.31 1.63
As shown in Table 4.1.3, when we separate the prompting strategies into Fact Only and Reduce Only, we observe that Fact Only still performs quite well, but Reduce Only leads to a significant decrease in performance. This suggests that while discouraging biased reasoning is important, providing explicit instructions that promote bias-free reasoning is crucial for mitigating bias in LLMs.
Moreover, comparing the performance of GPT-4 with GPT-3.5, we can see that the GPT-4’s improvements mostly stem from better (non-gender-related) world knowledge (making way less mistakes on factual questions) and self-consistency (less mistakes in the TF group).
More experimental results can be found in Appendix D.1, where we design prompts to control the degree to which we steer LLMs away from biased reasoning.
4.2 Demographic Bias: Discrim-Eval
@@
4.2.1 Experimental Settings (Discrim-Eval)
@@
To amplify bias-free reasoning (Fact), we first remove all demographic information in each scenario and ask the LLM to make a Yes or No decision based on this base scenario. Then we explicitly include its answer to the base scenario in the prompt and ask the model to decide on the original scenario which contains specific demographic information (e.g., age 20).
4.2.2 Our Results (Discrim-Eval)
@@
Each bar in the figure represents the relative gap, which measures the disparity in the probability of making an affirmative decision (‘Yes’) for the most and least privileged groups within each demographic category, characterizing the relative extent of discrimination. The bars are color-coded according to the methods used, where lighter shades correspond to the methods that explicitly encourage bias-free reasoning (e.g. purple bars denote Illegal and light purple bars denote Illegal + Fact). Default (black bar) denotes the prompt in the original scenario (with demographic information), without any modifications.
While different prompts of discouraging biased reasoning show various degrees of effectiveness in reducing demographic bias, no single prompt stands out as the universally most effective option across all LLMs and demographic categories. However, encouraging bias-free reasoning (Fact) collectively with discouraging biased reasoning can further decrease the relative gap, compared to employing the approaches on their alone. This observation is consistent across 4 LLMs and 3 demographic categories.
It is also noteworthy that none of the approaches achieves a zero gap. This indicates that while the debiasing prompt designs can reduce the discrimination in LLM outputs, they do not completely eliminate it.
Adverse Impact on Specific Demographic Groups

### v2

@@ 4.2.1 Experimental Settings (BBQ)
We use similar baselines as in WinoBias (Section 4.1.1), where we compare DDP with Default and the stronger baseline ICL with contrastive examples. We change the number of ICL examples to 8 to match the settings in Si et al. (2022).
Dual Directional Prompting on BBQ
is also similar to what we defined on WinoBias. To encourage fact-based reasoning, we follow Figure 1(b) to first convert an original question into a base question where we remove social category information by using neutral references (i.e., Person X and Person Y) to label the two entities in the given context. Then we ask the LLM the base question and prepend its answer before asking the original question (Strategy I). To discourage biased reasoning, we ask it not to use the information related to the underlying social category when making the decision (Strategy III).
4.2.2 Our Results (BBQ)
In Table 3, we can see that on common social categories where biases may already be well-regulated (e.g., age, gender, and race), even the Default method has a reasonably good performance. Yet, on social categories where bias may be more subtle (e.g., physical appearance, religion, socio-economic status, and sexual orientation), the model has a much lower accuracy when no additional prompts are used. On these categories both ICL and DDP significantly improved the performance over Default, and DDP achieves the highest performance across 8 out of 9 social categories.
Table 3: Performance comparison of various debiasing methods on BBQ. We measure how prompting-based debiasing methods help GPT-4 predict the correct choice given an adequately informative context.
Accuracy (%) $\uparrow$ Age Disability Status Gender Identity Nationality Physical Appearance Race Ethnicity Religion SES Sexual Orientation
Default 94.78 89.46 89.70 97.99 72.34 90.23 74.17 68.36 68.98
ICL (8-shot) 99.95 97.17 98.24 99.68 82.11 97.09 93.00 87.15 91.20
DDP (ours) 99.95 95.76 99.33 100.00 88.71 99.07 97.67 97.58 97.69
5 Conclusion
Our work presents a causality-guided, prompting-based debiasing framework for LLMs’ decisions. The framework tackles social biases and provides a novel perspective on how social category information can affect LLMs’ decisions via different causal pathways. Leveraging these causal insights, we propose principled, prompting-based debiasing strategies that utilize selection mechanisms to regulate the influence of bias. These strategies aim to discourage biased reasoning and encourage fact-based reasoning, unifying existing prompt-based debiasing methods and addressing their critical gap in promoting fact-based reasoning. Besides strong empirical results, we also prove that bias can be completely removed from LLMs’ decisions when the objectives in all three strategies are satisfied.
For future work, our proposed Dual Directional Prompting (DDP) method may also be used to identify pairs of positive (unbiased) and negative (biased) responses to learn a reward model (utilizing the intuition that an unbiased response should align with the model’s base decision).
Broader Impact
In this paper, we present a causality-based LLM output debiasing framework. We aim to provide causal understandings on both the underlying generating process for the training data corpus, and the LLM reasoning process where the output is modulated by input prompts through selection mechanisms. Our proposed framework serves as a general solution to promote fairness in pretrained LLMs without accessing their model parameters. This framework is applicable to many socially important domains, and we do not foresee any particular potential negative social impact or ethical concerns.
Acknowledgment
We would also like to acknowledge the support from NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Quris AI, Florin Court Capital, and MBZUAI-WIS Joint Program. ZT is supported by the National Institute of Justice (NIJ) Graduate Research Fellowship, Award No. 15PNIJ-24-GG-01565-RESS.
Appendix A Related Works
In this section, we provide detailed discussions on related works. We consider the combinations of very related topics including causality, algorithmic fairness, and LLM reasoning. In Section A.1, we reviewed existing literature on how they address selection. In Section A.2, we consider causal notions of fairness that do not specifically pertain to the LLM context. In Section A.3, we consider existing efforts to draw connections between causality and LLM reasoning. In Section A.4, we consider previous works on LLM debiasing. In Section A.5, we consider previous works that involve all three topics.
A.1 Selection mechanisms
Previous literature has investigated selection mechanisms from diverse perspectives. For instance, the influence of selection bias on statistical inference in economic and sociological studies (Heckman, 1979; 1990; Winship & Mare, 1992), causal discovery when there are selection variables and latent common causes (Spirtes et al., 1995; Zhang, 2008), the identification and estimation of functional causal models when selection exists (Zhang et al., 2016), the identifiability of causal effect in the presence of selection bias from the graphical condition perspective (Bareinboim & Pearl, 2012; Bareinboim & Tian, 2015; Correa et al., 2019) and from the potential outcome perspective (Hernán & Robins, 2020), and the identification of the existence of selection bias from observational data under certain functional assumptions (Kaltenpoth & Vreeken, 2023). Recent work also demonstrates the identification of selection structures in sequential data (Zheng et al., 2024).
A.2 Causality and Fairness
It has been recognized in the algorithmic fairness literature that causality provides a unique tool to facilitate a better understanding of the data-generating process, and therefore, more effective bias quantifications and mitigations (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019; Nabi et al., 2022; von Kügelgen et al., 2022; Tang et al., 2023a). Previous causal fairness literature has considered notions based on estimating or bounding various kinds of causal effects (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019), and also the causal modeling of the dynamics as well as the long-term implications of bias mitigation strategies (Creager et al., 2020; Zhang et al., 2020; Tang et al., 2023a; b).
A.3 Causality and LLMs
A.4 Debiasing Language Models
There is a large amount of work discussing bias and fairness in the context of language models (LMs) Bordia & Bowman (2019); Liang et al. (2021); Abid et al. (2021); Wang et al. (2023a); Liu et al. (2023); Ray (2023); Rozado (2023), and our investigation lies on debiasing techniques with causal understandings of the sources of biases. For debiasing approaches in the context of LMs, there are proposals involving direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), incorporating Reinforcement Learning with Human Feedback (RLHF) to better align the models with human values (Ouyang et al., 2022; Bai et al., 2022; Yao et al., 2023), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023; Furniturewala et al., 2024). We focus on prompting-based techniques and identify principles for prompt designs to steer LLMs toward unbiased responses by a). reducing biased reasoning and b). encouraging fact-based reasoning. We provide demonstrations of how we can employ the above two principles. Works on LLMs reasoning such as Wei et al. (2022); Zheng et al. (2023); Furniturewala et al. (2024) can also be incorporated into our framework to encourage fact-based reasoning(e.g., the Implication Prompting proposed in Furniturewala et al. (2024) can be viewed as an implicit approach encouraging bias-free reasoning). Though prompting-based debiasing methods become the only viable choice with black-box access, when we do have access to the model’s weights and enough compute resources to fine-tune the model, direct or instruction-based fine-tuning should be conducted first. Such approaches enable the integration of bias mitigation directly into the model parameters and can address the model’s intrinsic biases.
A.5 Debiasing LLMs from Causal Perspectives
Most closely related literature considers LLM debiasing strategies from causal perspectives. Vig et al. (2020) utilize causal mediation analysis and consider neuron-level intervention to investigate the instantiation of gender bias in Transformer-based language models (Vaswani et al., 2017). Zhang et al. (2024) also employs a similar view where they consider the chain-of-thought generated by LLMs as a mediator variable between prompt and answer. Wang et al. (2022) proposed a causal graph for relation extraction, and their counterfactual analysis lies in analyzing the change in the output relation distribution before and after removing the textual context. Zhou et al. (2023) propose Causal-Debias to mitigate the unwanted stereotypical association by fine-tuning pretrained language models. The causal model they consider involves four variables: label-relevant factor, bias-relevant factor, raw sentence, and ground-truth label. Wang et al. (2023b) pay special attention to entity bias and propose a specific structural causal model (SCM) for easier parameter estimations, such that the intervention-based mitigation strategy can be carried out. Their causal model involves four variables: entity, raw text, LLM input, and LLM output.
Appendix B Further Illustrations and Discussions of Our Framework
(c) LLM reasoning process (annotated with selection mechanisms specified by prompts)
(d) LLM reasoning process (Strategy I: nudge towards social-agnositic fact)
(e) LLM reasoning process (Strategy II: alternative fact counteracts existing selection bias)
(f) LLM reasoning process (Strategy III: nudge aware from social-salient text)
(g) LLM reasoning process (all strategies combined)
Figure 4: Additional illustrations on debiasing strategies.
In Figure 4, we use light coral (blue) to highlight unregulated (regulated) information flows from social category representations to LLM outputs in the internal reasoning process. We use mix-colored highlight to denote partially regulated information flow, e.g., the edge from social-salient text representation to LLM potential output in Figure 3(f). Comparing Figure 3(g) (all strategies combined) with Figures 3(d)3(f) (strategies applied individually), we can see that collectively, the strategies address the social bias in language more comprehensively.
Here by “regulated,” we are referring to the constraint over the association between social category information representation and the potential output. We would like to note that our causal models can naturally handle situations where certain kinds of neutral dependence patterns involving demographic information are not necessarily considered problematic (Tang et al., 2024). Recent benchmark suite by Wang et al. (2025) also explicitly argues that group difference awareness matters when measuring desired group discrimination in LLMs (e.g., in legal contexts).
We present relevant experimental results in Section D.1. We would also like to note that our causality-guided framework is not limited to situations where social category information is explicitly involved. One can adapt our framework to the specific practical setting by incorporating relevant nodes into causal graphs, thereby identifying the most suitable debiasing strategies therein.
Appendix C Experimental Details
Pretrained LLMs
For the GPT models used in our experiments on WinoBias and Discrim-Eval, we consider snapshots from June 13th, 2023 where the knowledge cut-off time is Sep 2021. Since the legacy GPT-3 model (a.k.a., text-davinci-003) is no longer supported when we conduct the experiments, we use the model “gpt-3.5-turbo-instruct” instead as it has similar capabilities as GPT-3 era models. The Mistral-7B model we use in our experiments is the improved instruction fine-tuned version (a.k.a., “Mistral-7B-Instruct-v0.2”). For Mistral-7B, we choose the instruction-finetuned version over the base version as the effectiveness of prompting strategies in regulating biased pathways is strongly tied to the model’s ability to follow instructions. We are unable to use Claude 2 for our experiments on the Discrim-Eval dataset because we cannot access the probabilities of generated tokens. All LLMs responses are obtained with a temperature of 0. For our experiments on BBQ data set, we use the latest GPT-4 version (i.e., gpt-4-turbo).
Iterative Prompting
C.1 Gender Bias: WinoBias
Since there are 571 examples where two or more models make the same mistakes on the factual questions. After looking into these mistakes, we find that some of the examples are indeed ambiguous in nature (e.g., "The construction worker stared at the teacher because [who] was angry"). Therefore, we collected human evaluations to understand the inherent ambiguity in the WinoBias dataset. For each of the 571 examples, we collect 5 human responses on the factual questions (i.e., which sentence is more likely to happen in real life), and we ask the human annotator to choose from {Sentence A, Sentence B, and Equally likely}. Based on the human responses, we identified 60 examples (55 from the Type I task and 5 from the Type II task) where three or more annotators disagree with the ground truth answers or think both sentences in the factual questions are equally likely to happen in real life. We removed these 60 examples during our evaluation, and we will include the human evaluation results in our codebase.
Also, for instances where the model was prompted more than three times but still failed to provide a definitive choice (e.g., when presented with options A and B but refused to select either), those refusal examples were removed from our calculations (e.g., in Tables 4.1.3C.2). This approach ensures the reported percentages accurately reflect the cases where the model made a conclusive decision.
The ICL baseline on the WinoBias dataset is from Si et al. (2022), where they constructed various sets of ICL examples (e.g., examples exclusively from anti-bias scenarios for Type I questions) and compared their effectiveness in debiasing LLMs’ outputs. Their findings reveal while using ICL examples from a specific category enhances performance for that category (e.g., pro-bias examples improve pro-bias predictions), a balanced set of pro-bias and anti-bias examples proved most effective in reducing bias. Therefore, this balanced strategy is one of the baselines we compare against.
C.2 Demographic Bias: Discrim-Eval
@@
Table 4: Error Analysis on WinoBias Type II coreference task. We divide the models’ responses into 4 categories to better understand their success and failure cases. TT denotes the (%) of examples where LLM answers both the factual question and the original question correctly, FF denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge. TF denotes the coreference errors caused by gender bias as only factual questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the factual questions get wrong but original questions are correctly answered.
GPT-4     DDP 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25 Fact Only 99.36 99.87 0.25 0.00 0.13 0.00 0.25 0.13 Counteract Only 98.98 99.49 0.76 0.13 0.00 0.38 0.25 0.00 Default 98.73 99.75 0.89 0.13 0.25 0.00 0.13 0.13
Table 5: Error analysis of Dual Directional Prompting  on Type I and Type II questions in WinoBias.
&82.47 82.34 1.27 1.40 0.76 1.91 15.50 14.36     GPT-3.5     Type I 70.15 78.70 10.58 2.04 2.58 5.97 16.55 13.16 Type II 88.69 89.71 1.14 0.13 3.43 4.70 6.73 5.46     Claude     Type I 73.00 73.00 2.99 2.99 1.09 2.17 22.93 21.85 Type II 79.80 81.19 6.61 5.21 2.54 4.07 11.05 9.53     GPT-4     Type I 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Type II 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25
Table 6: Ablation study employing Strategy III to reduce biased reasoning with Claude 2 on WinoBias
[…]
DDP (Strategy III) 55.81 83.59 27.78
DDP (Strategies II+III) 75.13 79.29 4.16
DDP (Strategies I+II) 74.08 75.17 1.09
DDP (Strategies I+II+III) 78.91 79.92 1.01
Table 7: Adjusting the levels of counteracting existing selection bias on GPT-3.5 with Dual Directional Prompting. We investigated how the level of counteracting stereotype impacts the model performance by adjusting the Fact part in the input prompt. For example, a counteract level of 100% indicates the following prompt: Assume that the physician can be male $0\%$ of the time and female $100\%$ of the time, and assume that the secretary can be male $0\%$ of the time and female $100\%$ of the time.
Appendix D Additional Experimental Results
D.1 Gender Bias: WinoBias
We include additional ablation studies on the WinoBias dataset, which include a detailed error analysis on WinoBias Type II coreference task (Table C.2), a detailed error analysis of Dual Directional Prompting with 4 LLMs on Type I and Type II tasks (Table C.2), an ablation study related to Strategy III—nudge aware from social-salient text—(Table 6), and an ablation study on adjusting the levels of biased reasoning (Table C.2) in the prompt design.
D.2 Demographic Bias: Discrim-Eval
@@
D.2.1 Experimental Settings (Discrim-Eval)
@@
To amplify fact-based reasoning (Fact), we first remove all demographic information in each scenario and ask the LLM to make a Yes or No decision based on this base scenario (Strategy I). Then we explicitly include its answer to the base scenario in the prompt and ask the model to decide on the original scenario which contains specific demographic information.
Figure 5: Performance comparison on Discrim-Eval across three demographics. The bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning). Encouraging fact-based reasoning universally decreases the relative gap when added with methods that reduce biased reasoning.
D.2.2 Our Results (Discrim-Eval)
@@
Each bar in the figure represents the relative gap, which measures the disparity in the probability of making an affirmative decision (‘Yes’) for the most and least privileged groups within each demographic category, characterizing the relative extent of discrimination. The bars are color-coded according to the methods used, where lighter shades correspond to the methods that explicitly encourage fact-based reasoning (e.g. purple bars denote Illegal and light purple bars denote Illegal + Fact). Default (black bar) denotes the prompt in the original scenario (with demographic information), without any modifications.
While different prompts of discouraging biased reasoning show various degrees of effectiveness in reducing demographic bias, no single prompt stands out as the universally most effective option across all LLMs and demographic categories. However, encouraging fact-based reasoning (Fact) collectively with discouraging biased reasoning can further decrease the relative gap, compared to employing the approaches on their alone. This observation is consistent across 4 LLMs and 3 demographic categories. It is also noteworthy that none of the approaches achieves a zero gap. This indicates that while the debiasing prompt designs can reduce the discrimination in LLM outputs, they do not completely eliminate it.
Adverse Impact on Specific Demographic Groups